OpenStudy (anonymous):

Three years ago Bob was twice as old as Adam will be four years from now. Adam is "a" years old now. What is Bob's age in terms of Adam's age?

OpenStudy (anonymous):

2a+4? I'm not 100% sure...

OpenStudy (anonymous):

2a+11

OpenStudy (anonymous):

b-3=2(a+4) b-3=2a+8 b=2a+11 b=bob a=adam